Klein Company distributes a high-quality bird feeder that sells for $50 per unit. Variable costs are $20 per unit, and fixed costs total $480,000 annually.

Required:

Answer the following independent questions:

1. What is the product’s CM ratio?

2. Use the CM ratio to determine the break-even point in sales dollars.

3. The company estimates that sales will increase by $65,000 during the coming year due to increased demand. By how much should operating income increase?

4. Assume that the operating results for last year were as follows:

  Sales $ 1,000,000
  Variable expenses 424,000
  Contribution margin 576,000
  Fixed expenses 480,000
  Operating income $ 96,000

a. Compute the degree of operating leverage at the current level of sales.

b. The president expects sales to increase by 14% next year. By how much should operating income increase?

5-a. Refer to the original data. Assume that the company sold 22,500 units last year. The sales manager is convinced that a 14% reduction in the selling price, combined with a $69,000 increase in advertising expenditures, would cause annual sales in units to increase by 30%. Prepare two contribution format income statements, one showing the results of last year’s operations and one showing what the results of operations would be if these changes were made. (Round "Per Unit" answers to 2 decimal places.)

5-b. Would you recommend that the company do as the sales manager suggests?

  • Yes

  • No

6. Refer to the original data. Assume again that the company sold 22,500 units last year. The president feels that it would be unwise to change the selling price. Instead, he wants to increase the sales commission by $2 per unit. He thinks that this move, combined with some increase in advertising, would increase annual unit sales by 50%. By how much could advertising be increased with profits remaining unchanged? Do not prepare an income statement; use the incremental analysis approach.

hi chegg team can you please solve this question for me...
Answer #1

1) CM ratio = (50-20)/50 = 60%

2) Break even sales = 480000/.60 = $800000

3) Net operating income increase by 65000*60% = 39000

4a) Degree of operating leverage = Contribution margin/Operating income = 576000/96000 = 6 Times

4b) Net operating income increase by 14*6 = 84%

Note : Please post whole question in the group of 4-4 sub parts as per Chegg guidelines